OpenStudy (ryanchick):

pleas help me with 255 over 48

Mehek (mehek14):

\(\dfrac{255}{48}\) ?

OpenStudy (ryanchick):

ya

Mehek (mehek14):

you can use a calculator

OpenStudy (ryanchick):

i did but it needs to become a fraction

Mehek (mehek14):

oh oke how many times does 48 go into 255 as a whole number?

OpenStudy (ryanchick):

5.3125 idk

Mehek (mehek14):

so 5 times what is 48 * 5?

OpenStudy (ryanchick):

240

Mehek (mehek14):

255 - 240?

OpenStudy (ryanchick):

15

Mehek (mehek14):

\(5\dfrac{15}{48}\) simplify 15/48

OpenStudy (ryanchick):

how do you do that

Mehek (mehek14):

what number goes in 15 and 48

OpenStudy (ryanchick):

3

Mehek (mehek14):

so 15/3 and 48/3

OpenStudy (mathstudent55):

If you reduce first, it makes the numbers easier to work with. \(\dfrac{255}{48} = \dfrac{3 \times 85}{3 \times 16} = \dfrac{85}{16} = \dfrac{5 \times 16 + 5}{16} = \dfrac{5 \times 16}{16} + \dfrac{5}{16} = 5 \dfrac{5}{16}\)

OpenStudy (ryanchick):

thank you